OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1

OpenStudy (anonymous):

-5x - 1 because y = mx+c

OpenStudy (cwrw238):

general for is y = mx + c